OpenStudy (anonymous):

what is the value of the y variable in the solution to the following system of equations. 3x-6y=12 -2x+3y=6 answers are -14 24 14 -24

OpenStudy (anonymous):

14.

OpenStudy (anonymous):

How is it 14?

OpenStudy (anonymous):

sorry, -14.

OpenStudy (anonymous):

how is it -14?

OpenStudy (anonymous):

eliminate x from the above.

OpenStudy (anonymous):

I don't get it?

OpenStudy (jdoe0001):

3x-6y=12 -2x+3y=6 in this case you'd want to know what's "y" so, you eliminate the other variable, "x", so "y" is isolated you change the top and/or bottom equation by usually multiplying for something that'd give you 2 opposites for the variable you'd want eliminated so if you had 3x-6y=12 how to eliminate "3x"? well "3x - 3x =0", so with a -3x something like 3x-6y=12 -3x -whatever gives 3x-6y=12 -3x-whatever -------------- 0 - whatever

OpenStudy (jdoe0001):

lemme do yours

OpenStudy (anonymous):

that you will get -14 right

OpenStudy (jdoe0001):

3x -6y = 12 times 2 -2x +3y = 6 times 3 $$ \begin{matrix} \color{red}{6x} &-12y & = &24 & \times 2\\ \color{red}{-6x} &+9y & = & 18 &\times 3\\ \hline\\ 0 & -3y & = & 42 \end{matrix} $$

OpenStudy (jdoe0001):

see the elimination?

OpenStudy (anonymous):

yea but what will the answer be then

OpenStudy (jdoe0001):

$$ \begin{matrix} \color{red}{6x} &-12y & = &24 & \times 2\\ \color{red}{-6x} &+9y & = & 18 &\times 3\\ \hline\\ 0 & -3y & = & 42 \end{matrix}\\ \\ \color{blue}{-3y = 42} \implies \color{blue}{y = \cfrac{42}{-3}} $$

OpenStudy (anonymous):

I see it that help me so much thank you

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10 2 -1 1 -2 can you help with this one too lol

OpenStudy (jdoe0001):

well, you want "x", so eliminate "y" off 2x + 3y = 1 3x - 4y = 10

OpenStudy (anonymous):

3x-6y=12 3x+6*2-x=4 3x+12-x=4 2x+12=4 x+12=4/2 x=2-12 x=-14

OpenStudy (anonymous):

what will you do next?

OpenStudy (anonymous):

working on it...lol

OpenStudy (jdoe0001):

heheh

OpenStudy (anonymous):

vaibhavm thank you lol

OpenStudy (anonymous):

well, you want "x", so eliminate "y" off 2x + 3y = 1 3x - 4y = 10 what do we do next lol. I am really bad at math.

OpenStudy (jdoe0001):

well, we take a peek at the "y's"

OpenStudy (jdoe0001):

2x + 3y = 1 3x - 4y = 10 hmmmm, we have +3y and -4y so, how can we get those 2 guys to be the same number, different sign, that is if one is 12 the other is -12 12 -12 =0, so eliminated

OpenStudy (anonymous):

I get that.

OpenStudy (anonymous):

what do we need to do next

OpenStudy (jdoe0001):

so, what number would you multiply +3x to get a \(\color{blue}{\text{number in positive?}}\) and what number would you multiply -4y to get the \(\color{red}{\text{number in negative?}}\)

OpenStudy (anonymous):

I don't know :(

OpenStudy (jdoe0001):

$$ \begin{matrix} 2x& + 3y &= &1 & \times 4 \\ 3x &- 4y &= &10 & \times 3\\ \end{matrix} \implies \begin{matrix} 8x& \color{red}{+ 12y} &= &4 & \\ 9x &\color{red}{- 12y} &= &30 &\\ \hline\\ ? & ? & & ? \end{matrix}\\ $$

OpenStudy (jdoe0001):

so, what values would that give you?

OpenStudy (anonymous):

I think it would be the 8x i don't know but +12y is 0 and 4 I don't know sorry 9x -12y 30

OpenStudy (jdoe0001):

well, you ADD THEM vertically

OpenStudy (anonymous):

so it be 17 0 34 right

OpenStudy (jdoe0001):

17x + 0 = 34, yes, now just solve for "x" :)

OpenStudy (jdoe0001):

or 17x = 34

OpenStudy (anonymous):

what happened to your first question.....did we get an answer?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

yes lol its -14

OpenStudy (anonymous):

coool...... :)

OpenStudy (anonymous):

17x + 0 = 34, yes, now just solve for "x" :) or 17x = 34 how do i solve that lol.

OpenStudy (anonymous):

vaibhavm yups lol

OpenStudy (anonymous):

X=2

OpenStudy (anonymous):

how did you get 2

OpenStudy (anonymous):

17x=34 x=34/17 x=2

OpenStudy (jdoe0001):

ahem you divide both sides by 17

OpenStudy (anonymous):

I see now. thank you

OpenStudy (jdoe0001):

$$ 17x =34 \implies \cfrac{17x}{17} = \cfrac{34}{17} $$

OpenStudy (anonymous):

lol.....you are welcome.....you can give me a medal......lol

OpenStudy (anonymous):

lol but i need more help lol

OpenStudy (anonymous):

lol....i m there....pleasure.

OpenStudy (anonymous):

The graph below is represented by which system of inequalities? -x + 3y < 9 y > -2x + 1 -x + 3y 9 y > -2x + 1 -x + 3y < 9 y -2x + 1 None of these systems represent the graph shown. see how do i do this lol

OpenStudy (anonymous):

Lol ok ;)

OpenStudy (anonymous):

lol........ i quit.....hahahahaha....lol

OpenStudy (anonymous):

Nooooooooooooooooooooooooooo lol why you don't want to help me :(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

how do i do this The graph below is represented by which system of inequalities? -x + 3y < 9 y > -2x + 1 -x + 3y 9 y > -2x + 1 -x + 3y < 9 y -2x + 1 None of these systems represent the graph shown.

OpenStudy (anonymous):

lol......its too complicated....lol

OpenStudy (anonymous):

lol yea it is

OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12 3y + 12 -3y + 12 3y - 12 -3y - 12 do you know how do this lol